# A Wasserstein perspective of Vanilla GANs

### v1

@@ 1 Introduction
to learn an unknown distribution $\mathbb{P}^{*}$ of the random variable $X$. The generator $G$ chosen from a set $\mathcal{G}$ applied to the latent random variable $Z$ aims to mimic the distribution of $X$ as closely as possible. The discriminator $D,$ chosen from a set $\mathcal{D},$ has to distinguish between real and fake samples.
@@
are obtained by replacing the Jensen-Shannon divergence by the Kantorovich dual of the Wasserstein distance. Here, $\operatorname*{Lip}(1)$ denotes the set of all Lipschitz continuous functions with Lipschitz constant bounded by one. This approach can be generalized to Integral Probability Metrics (IPM, Mueller, (1997)) GANs, see Liang, (2021).
The analysis of Wasserstein GANs can exploit the existing theory on the Wasserstein distance. The latter has a long record of research, particularly in the context of optimal transport (Villani,, 2008) but also in machine learning, see Torres et al., (2021) for an overview. In contrast, Vanilla GANs and the Jensen-Shannon divergence have been studied less extensively, and fundamental questions have not been settled. In particular, all statistical results for Vanilla GANs require the same dimension of the latent space and the target space which is in stark contrast to common practice. Another algorithmic drawback of Vanilla GANs highlighted by Arjovsky & Bottou, (2017) is that an arbitrarily large discriminator class prevents the generator from learning. The Jensen-Shannon divergence between singular measures is by definition maximal. Therefore, we cannot expect convergence in a dimension reduction setting. In practice, however, Vanilla GANs have worked in various settings. Thus, using neural networks as a discriminator class must be advantageous compared to the set of all measurable functions. This empirical fact is supported by the numerical results by Farnia & Tse, (2018) who impose a Lipschitz constraint on the discriminator class.
@@
Our work aims to bridge the gap between Vanilla GANs and Wasserstein GANs while addressing the theoretical limitations of the former ones. By imposing a Lipschitz condition on the discriminator class in (1), we recover a Wasserstein-like behavior. As a main result, we can derive an oracle inequality for the Wasserstein distance between the true data generating distribution and its Vanilla GAN estimate. In particular, this allows us to transfer key features, such as dimension reduction, known from the statistical analysis of Wasserstein GANs. We show that the statistical error of the modified Vanilla GAN depends only on the dimension of the latent space, independent of the potentially much larger dimension of the feature space $\mathcal{X}$. Thus Vanilla GANs can avoid the curse of dimensionality. Such properties are well known from practice, but cannot be verified by the classical Jensen-Shannon analysis. On the other hand the derived rate of convergence for the Vanilla GAN is slower than for Wasserstein GANs which is in line with the empirical advantage of Wasserstein GANs.
@@
The existence and uniqueness of the optimal generator for Vanilla GANs is shown by Biau et al., (2018) under the condition that the class $\mathcal{G}$ is convex and compact. They also study the asymptotic properties of Vanilla GANs. Belomestny et al., (2021) have shown a non-asymptotic rate of convergence in the Jensen-Shannon divergence for Vanilla GANs with neural networks under the assumption that the density of $\mathbb{P}^{*}$ exists and that the generator functions are continuously differentiable.
In practice, however, the Rectifier Linear Unit activation function (ReLU activation function) is commonly used (Aggarwal,, 2018, p.13). The resulting neural network generates continuous piecewise linear functions. Therefore, the convergence rate of Belomestny et al., (2021) combined with Belomestny et al., (2023) is not applicable to this class of functions.
@@
Convergence rates with respect to the Wasserstein distance have been studied by Chae, (2022) and Chen et al., (2020). Up to a logarithmic factor, optimal rates in the Hölder metric were obtained by Stéphanovitch et al., (2023) using smooth networks. In a similar setting, Chakraborty & Bartlett, (2024) discussed several methods for dimension reduction. Recently, Suh & Cheng, (2024) have reviewed the theoretical advances in Wasserstein GANs.
@@
(Khromov & Singh, (2023), in context of GANs see Than & Vu, (2021)). Implementations of Wasserstein GANs have evolved from weight clipping (Arjovsky et al.,, 2017) to penalizing the objective function (Gulrajani et al.,, 2017; Wei et al.,, 2018; Zhou et al.,, 2019; Petzka et al.,, 2017; Miyato et al.,, 2018; Asokan & Seelamantula,, 2023), which heuristically leads to networks with bounded Lipschitz constants. Farnia & Tse, (2018) use an objective function that combines Wasserstein and Vanilla GANs.
@@
In Section 2 we introduce the Vanilla GAN distance, which characterizes the optimization problem (1). In Section 3, we investigate the relation between the Vanilla GAN distance and the Wasserstein distance. We show that the distances are compatible to each other while not being equivalent. Using this relation, we derive an oracle inequality for the Vanilla GAN in Section 4, where $\mathcal{G}$ is an arbitrary set and $\mathcal{D}$ is a set of Lipschitz functions. We show that Vanilla GANs can avoid the curse of dimensionality. In Section 5 we consider the situation where $\mathcal{G}$ and $\mathcal{D}$ consist of neural networks. Here we relax the Lipschitz condition to a $\alpha$-Hölder condition and prove a quantitative result for the approximation of a Lipschitz function by a feedforward ReLU network with bounded Hölder norm. We then prove a convergence rate for the Vanilla GAN with network generator and discriminator. All proofs are deferred to section 7.
@@ 2 The Vanilla GAN distance
Let us first fix some notation. We equip $\mathbb{R}^{d}$ with the $\ell_{p}$-norm $|x|_{p}$, $1\leq p\leq\infty$, and denote the number of nonzero entries of a $k\times l$ matrix $A,$ where $k,l\in\mathbb{N},$ by $|A|_{\ell^{0}}\coloneqq|\{(i,j)\colon A_{ij}\neq 0\}|.$ For ease of notation we abbreviate for $x\in(0,\infty)$
@@
$\operatorname*{Lip}(L,B,\Omega)\coloneqq\{f\colon\Omega\rightarrow\mathbb{R}\mid\|f\|_{\infty,\Omega}\leq B,\ |f(x)-f(y)|\leq L|x-y|_{p},\ \forall x,y\in\Omega\}.$
The set of unbounded Lipschitz functions is denoted by $\operatorname*{Lip}(L,\Omega)\coloneqq\operatorname*{Lip}(L,\infty,\Omega).$ By Rademacher’s theorem, a Lipschitz function is differentiable almost everywhere. For $\alpha\in(0,1]$ we define the $\alpha$-Hölder norm by
$\|f\|_{\mathcal{H}^{\alpha}(\Omega)}:=\max\left\{\|f\|_{{\infty}},\;\underset{x,y\in\Omega}{\operatorname{ess\;sup}}\;\frac{|f(x)-f(y)|}{|x-y|_{p}^{\alpha}}\right\}$ (4)
@@
We observe an i.i.d. sample $X_{1},...,X_{n}\sim\mathbb{P}^{*}$ with values in $\mathcal{X}\coloneqq(0,1)^{d}$. On another space $\mathcal{Z}\coloneqq(0,1)^{d^{*}}$, called the latent space, we choose a latent distribution $\mathbb{U}$. Unless otherwise specified, $X\sim\mathbb{P}^{*}$ and $Z\sim\mathbb{U}$. Throughout, the generator class $\mathcal{G}$ is a compact nonempty set of measurable functions from $\mathcal{Z}$ to $\mathcal{X}$. For $G\in\mathcal{G}$ the distribution of the random variable $G(Z)$ is denoted by $\mathbb{P}^{G(Z)}.$
@@
in terms of the Vanilla GAN distance between probability measures $\mathbb{P}$ and $\mathbb{Q}$
@@
To choose the generator $\hat{G}_{n}$ as the empirical risk minimizer, the unknown distribution $\mathbb{P}^{*}$ in (6) must be replaced by the empirical distribution $\mathbb{P}_{n}$ based on the observations $X_{1},...,X_{n}$. In practice, the expectation with respect to $Z\sim\mathbb{U}$ is replaced by an empirical mean too, which we omit for the sake of simplicity. Following Huang et al., (2022), the next and all subsequent results easily extend to the corresponding setting.
@@
Assume that $\mathcal{G}$ is chosen such that a minimum exists. Let $\mathcal{W}$ be symmetric, that is, for $W\in\mathcal{W},-W\in\mathcal{W}.$ For
@@
If $\mathcal{W}$ is the set of all measurable functions, the analysis by Goodfellow et al., (2014, Theorem 1) shows that the Vanilla GAN distance is equivalent to the Jensen-Shannon distance shifted by $\log(4)$. Arjovsky & Bottou, (2017) have elaborated on the theoretical and practical disadvantages of this case. Similar to the total variation distance or the Hellinger distance, the Jensen-Shannon divergence is not compatible with high-dimensional settings because it cannot distinguish between different singular measures. Therefore, we need a weaker distance and thus restrict $\mathcal{W}$.
The key insight of Wasserstein GANs (2) is that this particular drawback of the Jensen-Shannon distance is can be solved by the Wasserstein distance. The latter is a metric on the space of probability distributions with finite first moment and meterizes weak convergence in this space (Villani,, 2008, Theorem 6.9). Let $\mathbb{P}$ and $\mathbb{Q}$ be two probability distributions on the same measurable space $(\Omega,\mathcal{A})$, the Wasserstein-$1$ distance is defined as
@@ 3 From Vanilla to Wasserstein and back
For $L>2$ and $B>0$ we have
$\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,B^{\prime})\\ W(\cdot)>-\log(2-2/L)\end{subarray}}$$\Big{\{}\mathbb{E}[W(X)-W(Y)]-\frac{L(L-1)}{2}\mathbb{E}[W(X)^{2}]\Big{\}}$
$\leq\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})\leq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>-\log(2)\end{subarray}}\Big{\{}\mathbb{E}[W(X)-W(Y)]-\frac{e^{B}}{(2e^{B}-1)^{2}}\mathbb{E}[W(X)^{2}]\Big{\}},$
@@
Theorem 3.1 reveals that the Vanilla GAN distance is indeed compatible with the Wasserstein distance and will allow us to prove rates of convergence of the Vanilla GAN with respect to the Wasserstein distance. In doing so, we need to consider the consequences of the penalty term. More precisely, we can deduce from Theorem 3.1 that the Vanilla GAN distance is bounded from above and below by the Wasserstein distance or the squared Wasserstein distance, respectively.
@@
Let $L>2$ and $B\in(0,\infty].$ For two probability measures $\mathbb{P}$ and $\mathbb{Q}$ on $\mathcal{X}$ we have
@@
The assumption that $L>2$ is not very restrictive and takes into account that in practically relevant cases, such as neural network discriminators, the Lipschitz constant is typically quite large. More importantly, we observe a gap between $\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})^{2}$ in the lower bound and $\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})$ in upper bound when $\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})<1$ which is a consequence of the penalty term in Theorem 3.1. The following example indicates that this loss is unavoidable, by restricting the discriminator class to a subset of $\operatorname*{Lip}(L)$.
@@
See Appendix A for more details on these calculations.
Wasserstein GANs, where the generator is chosen as the empirical risk minimizer of the Wasserstein distance (18), achieve optimal convergence rates (up to logarithmic factors) with respect to the Wasserstein distance as proved by Stéphanovitch et al., (2023). In view of Theorem 3.2 we cannot hope that Vanilla GANs achieve the same rate even if we use a Lipschitz discriminator class. This is in line with the better performance of Wasserstein GANs in practice. However, Theorem 3.2 allows us to study the behavior of Vanilla GANs in settings where the dimension of the latent space is smaller than the dimension of the sample space, a setting that is excluded in all previous work on convergence rates for Vanilla GANs.
4 Oracle inequality for Vanilla GAN in Wasserstein distance
Our aim is to bound the Wasserstein distance between the unknown distribution $\mathbb{P}^{*}$ and the generated distribution $\mathbb{P}^{\hat{G}_{n}(Z)}$ using the empirical risk minimizer $\hat{G}_{n}$ of the Vanilla GAN. The following oracle inequality shows that imposing a Lipschitz constraint on the discriminator class does circumvents the theoretical limitations of Vanilla GANs which is caused by the coupling with the Jensen-Shannon distance. Recall notation (3).

### v2

@@ 1 Introduction
to learn an unknown distribution $\mathbb{P}^{*}$ of the random variable $X$. The generator $G$ chosen from a set $\mathcal{G},$ applied to the latent random variable $Z$ aims to mimic the distribution of $X$ as closely as possible. The discriminator $D,$ chosen from a set $\mathcal{D},$ has to distinguish between real and fake samples.
@@
are obtained by replacing the Jensen-Shannon divergence by the Kantorovich dual of the Wasserstein distance. Here, $\operatorname*{Lip}(1)$ denotes the set of all Lipschitz continuous functions with a Lipschitz constant bounded by one. This approach can be generalized using Integral Probability Metrics (IPMs, Mueller, (1997)). For the application to GANs, see for example Liang, (2021).
The analysis of Wasserstein GANs can exploit the existing theory on the Wasserstein distance. The latter has a long record of research, particularly in the context of optimal transport (Villani,, 2008) but also in machine learning, see Torres et al., (2021) for an overview. In contrast, Vanilla GANs and the Jensen-Shannon divergence have been studied less extensively, and fundamental questions have not been settled. In particular, all statistical results for Vanilla GANs require the same dimension of the latent space and the target space which is in stark contrast to common practice. Another algorithmic drawback of Vanilla GANs highlighted by Arjovsky & Bottou, (2017) is that an arbitrarily large discriminator class prevents the generator from learning. The Jensen-Shannon divergence between singular measures is by definition maximal. Therefore, we cannot expect proofs of convergence in a dimension reduction setting. In practice, however, Vanilla GANs have worked in various settings. Thus, using neural networks as a discriminator class must be advantageous compared to the set of all measurable functions. This empirical fact is supported by the numerical results by Farnia & Tse, (2018) who impose a Lipschitz constraint on the discriminator class.
A wide range of methods exists for measuring distances between probability measures with varying properties. In view of the manifold hypothesis for high-dimensional data, it is crucial that the selected metric can discriminate between different singular measures. This is not the case for the Total Variation distance, the Jensen-Shannon distance, or even stronger metrics where singular distributions always attain the maximal value of the distance. Conversely, it is advantageous to select a strong metric, as this yields immediate bounds in weaker metrics. In view of this tension we will analyze Vanilla GANs with respect to the Wasserstein-$1$ distance which is often used in the statistical as well as the machine learning literature. A comprehensive overview of the advantages of the Wasserstein-$1$ distance over other measures that metrize weak convergence, we direct the reader to (Villani,, 2008, p. 98 f.). For an overview of distances between probability measures, we refer the reader to Gibbs & Su, (2002, Figure 1).
@@
Our work aims to bridge the gap in theoretical analysis between Vanilla GANs and Wasserstein GANs while addressing the theoretical limitations of the former ones. By imposing a Lipschitz condition on the discriminator class in (1), we recover Wasserstein GAN-like behavior. As a main result, we can derive an oracle inequality for the Wasserstein distance between the true data generating distribution and its Vanilla GAN estimate. In particular, this allows us to transfer key features, such as dimension reduction, known from the statistical analysis of Wasserstein GANs. We show that the statistical error of the modified Vanilla GAN depends only on the dimension of the latent space, independent of the potentially much larger dimension of the feature space $\mathcal{X}$. Thus Vanilla GANs can avoid the curse of dimensionality. Such properties are well known from practice, but cannot be verified by the classical Jensen-Shannon analysis. On the other hand the derived rate of convergence for the Vanilla GAN is slower than for Wasserstein GANs which is in line with the empirical advantage of Wasserstein GANs.
@@
We use a simple illustrative example to assess the practical implications of our theoretical results. This example allows us to quantify the rate depending on the number of observations, the dimension reduction property, and the stabilizing effect of a Lipschitz-constrained discriminator class.
@@
The existence and uniqueness of the optimal generator for Vanilla GANs is shown by Biau et al., (2020) under the condition that the class $\mathcal{G}$ is convex and compact. They also study the asymptotic properties of Vanilla GANs. Puchkin et al., (2024) have shown a non-asymptotic rate of convergence in the Jensen-Shannon divergence for Vanilla GANs with neural networks under the assumption that the density of $\mathbb{P}^{*}$ exists and that the generator functions are continuously differentiable.
In practice, however, the Rectifier Linear Unit activation function (ReLU activation function) is commonly used (Aggarwal,, 2018, p.13). The resulting neural network generates continuous piecewise linear functions. Therefore, the convergence rate of Puchkin et al., (2024) combined with Belomestny et al., (2023) is not applicable to this class of functions.
@@
Convergence rates with respect to the Wasserstein distance have been studied by Chen et al., (2020) and Chae, (2022). Up to a logarithmic factor, optimal rates in the Hölder metric were obtained by Stéphanovitch et al., (2023) using smooth networks. In a similar setting, Chakraborty & Bartlett, (2024) discussed several methods for dimension reduction. Recently, Suh & Cheng, (2024) have reviewed the theoretical advances in Wasserstein GANs.
@@
(Khromov & Singh, (2024), in context of GANs see Than & Vu, (2021)). Implementations of the Lipschitz constrained discriminator have evolved from weight clipping (Arjovsky et al.,, 2017) to penalizing the objective function (Gulrajani et al.,, 2017; Wei et al.,, 2018; Zhou et al.,, 2019; Petzka et al.,, 2018; Miyato et al.,, 2018; Asokan & Seelamantula,, 2023), which heuristically leads to networks with bounded Lipschitz constants. Farnia & Tse, (2018) use an objective function that combines Wasserstein and Vanilla GANs.
@@
In Section 2 we introduce the Vanilla GAN distance, which characterizes the optimization problem (1). In Section 3, we investigate the relation between the Vanilla GAN distance and the Wasserstein distance. We show that the distances are compatible to each other while not being equivalent. Using this relation, we derive an oracle inequality for the Vanilla GAN in Section 4, where $\mathcal{G}$ is a nonempty compact set and $\mathcal{D}$ is a set of Lipschitz functions. We show that Vanilla GANs can avoid the curse of dimensionality. In Section 5 we consider the situation where $\mathcal{G}$ and $\mathcal{D}$ consist of neural networks. Here we relax the Lipschitz condition to a $\alpha$-Hölder condition and prove a quantitative result for the approximation of a Lipschitz function by a feedforward ReLU network with bounded Hölder norm. We then prove a convergence rate for the Vanilla GAN with network generator and discriminator. In Section 6 we obtain a convergence rate for Wasserstein-type GANs with network generator and discriminator using our approximation result. This enables us to compare Vanilla GANs directly to Wasserstein GANs. In Section 7 we illustrate our theoretical results in a numerical example based on synthetic data. All proofs are deferred to Appendix A.
@@ 2 The Vanilla GAN distance
Let us first fix some notation. We equip $\mathbb{R}^{d}$ with the $\ell_{p}$-norm $|x|_{p}$, $1\leq p\leq\infty$, denote the number of nonzero entries of a $k\times l$ matrix $A,$ where $k,l\in\mathbb{N},$ by $|A|_{\ell^{0}}\coloneqq|\{(i,j)\colon A_{ij}\neq 0\}|,$ and define the ceiling of $x\in\mathbb{R}$ as $\lceil x\rceil:=\min\{k\in\mathbb{Z}\mid k\geq x\}.$ For ease of notation we abbreviate for $x\in(0,\infty)$
@@
$\operatorname*{Lip}(L,B,\Omega)\coloneqq\Big{\{}f\colon\Omega\rightarrow\mathbb{R}\,\Big{|}\,\|f\|_{\infty,\Omega}\leq B,\ \;\frac{|f(x)-f(y)|}{|x-y|_{p}}\leq L,x,y\in\Omega\Big{\}}.$
The set of unbounded Lipschitz functions is abbreviated by $\operatorname*{Lip}(L,\Omega)\coloneqq\operatorname*{Lip}(L,\infty,\Omega).$ By Rademacher’s theorem, a Lipschitz function is differentiable almost everywhere. For $\alpha\in(0,1]$ we define the $\alpha$-Hölder norm by
$\|f\|_{\mathcal{H}^{\alpha}(\Omega)}:=\max\Big{\{}\|f\|_{{\infty}},\;\underset{x,y\in\Omega}{\operatorname{ess\;sup}}\;\frac{|f(x)-f(y)|}{|x-y|_{p}^{\alpha}}\Big{\}}$ (4)
@@
We observe i.i.d. samples $X_{1},...,X_{n}\sim\mathbb{P}^{*}$ with values in $\mathcal{X}\coloneqq(0,1)^{d}$. On another space $\mathcal{Z}\coloneqq(0,1)^{d^{*}}$, called the latent space, we choose a latent distribution $\mathbb{U}$. Unless otherwise specified, $X\sim\mathbb{P}^{*}$ and $Z\sim\mathbb{U}$. We further assume that $\mathbb{P}^{*}$ and $\mathbb{U}$ have finite first moments. Throughout, the generator class $\mathcal{G}$ is a nonempty set of measurable functions from $\mathcal{Z}$ to $\mathcal{X}$. For $G\in\mathcal{G}$ the distribution of the random variable $G(Z)$ is denoted by $\mathbb{P}^{G(Z)}.$
@@
in terms of the Vanilla GAN distance between probability measures $\mathbb{P}$ and $\mathbb{Q}$ on $\mathcal{X}$
@@
To choose the generator $\hat{G}_{n}$ as the empirical risk minimizer, the unknown distribution $\mathbb{P}^{*}$ in (6) must be replaced by the empirical distribution $\mathbb{P}_{n}$ based on the observations $X_{1},...,X_{n}$. In practice, the expectation with respect to $Z\sim\mathbb{U}$ is replaced by an empirical mean too, which we omit for the sake of simplicity. Along Huang et al., (2022), the next and all subsequent results easily extend to the corresponding setting.
@@
Assume that $\mathcal{G}$ is chosen such that a minimum exists. Let $\mathcal{W}$ be symmetric, that is, $W\in\mathcal{W}$ implies $-W\in\mathcal{W}.$ For
@@
If $\mathcal{W}$ is the set of all measurable functions, the analysis by Goodfellow et al., (2014, Theorem 1) shows that the Vanilla GAN distance is equivalent to the Jensen-Shannon distance. Arjovsky & Bottou, (2017) have elaborated on the theoretical and practical disadvantages of this case. Similar to the Total Variation distance or the Hellinger distance, the Jensen-Shannon divergence is not compatible with high-dimensional settings because it cannot distinguish between different singular measures. Therefore, we need a weaker distance and thus restrict $\mathcal{W}$.
The key insight of Wasserstein GANs (2) is that this particular drawback of the Jensen-Shannon distance can be solved by the Wasserstein distance. The latter is a metric on the space of probability distributions with finite first moment and meterizes weak convergence in this space (Villani,, 2008, Theorem 6.9). Let $\mathbb{P}$ and $\mathbb{Q}$ be two probability distributions on the same measurable space $(\Omega,\mathcal{A})$, the Wasserstein-$1$ distance is defined as
@@
Hence, we throughout assume the following:
Assumption 1.
$\mathcal{G}$ is compact with respect to the supremum norm.
In the context of neural networks the compactness assumption is satisfied for all practically relevant implementations. Furthermore, it should be noted that the aforementioned assumption is only required for the use of the minimizing argument.
@@ 3 From Vanilla to Wasserstein and back
For $L>2$ and $B>0$ we have for probability measures $\mathbb{P}$ and $\mathbb{Q}$ on $\mathcal{X}$
$\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(1,B^{\prime})\\ W(\cdot)>-\log(2-2/L)\end{subarray}}$$\Big{\{}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}[W(X)-W(Y)]-\frac{L(L-1)}{2}\mathbb{E}_{X\sim\mathbb{Q}}[W(X)^{2}]\Big{\}}$
$\leq\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})\leq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>-\log(2)\end{subarray}}\Big{\{}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}[W(X)-W(Y)]-\frac{e^{B}}{(2e^{B}-1)^{2}}\mathbb{E}_{X\sim\mathbb{Q}}[W(X)^{2}]\Big{\}},$
@@
Theorem 3.1 reveals that the Vanilla GAN distance is indeed compatible with the Wasserstein distance and will allow us to prove rates of convergence of the Vanilla GAN with respect to the Wasserstein distance. In doing so, we need to investigate the consequences of the penalty term. An upper bound without the penalty term and independent of $B$ can be shown as in the proof of Theorem 3.2. For the lower bound, a similiar improvement cannot be expected in general as indicated in Example 3.3. However, the restriction to $\operatorname{Lip}(1,B^{\prime})$ has far less severe consequences than the corresponding restriction in the upper bound.
We can deduce from Theorem 3.1 that the Vanilla GAN distance is bounded from above and below by the Wasserstein distance or the squared Wasserstein distance, respectively.
@@
[…]
@@
The assumption $L>2$ is not very restrictive. In practically relevant cases, such as neural network discriminators, the Lipschitz constant is typically quite large. A higher Lipschitz constraint on the discriminator will subsequently result in a less stringent constraint on the neural network. However, an arbitrarily large Lipschitz constant is also undesirable, as the upper bound grows linearly in $L$.
More importantly, we observe a gap between $\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})^{2}$ in the lower bound and $\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})$ in upper bound when $\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})<1$ which is a consequence of the penalty term in Theorem 3.1. The following example indicates that this loss is unavoidable, by restricting the discriminator class to a subset of $\operatorname*{Lip}(L)$.
@@
See Section A.7 for more details on these calculations.
Wasserstein GANs, where the generator is chosen as the empirical risk minimizer of the Wasserstein distance (10), achieve optimal convergence rates (up to logarithmic factors) with respect to the Wasserstein distance as proved by Stéphanovitch et al., (2023), see also Section 6. In view of Theorem 3.2 we cannot hope that Vanilla GANs achieve the same rate even if we use a Lipschitz discriminator class. This is in line with the better performance of Wasserstein GANs in practice. However, Theorem 3.2 allows us to study the behavior of Vanilla GANs in settings where the dimension of the latent space is smaller than the dimension of the sample space, a setting that is excluded in all previous works on convergence rates for Vanilla GANs.
4 Oracle inequality for Vanilla GANs in Wasserstein distance
Our aim is to bound the Wasserstein distance between the unknown distribution $\mathbb{P}^{*}$ and the generated distribution $\mathbb{P}^{\hat{G}_{n}(Z)}$ using the empirical risk minimizer $\hat{G}_{n}$ of the Vanilla GAN. The following oracle inequality shows that imposing a Lipschitz constraint on the discriminator class does circumvent the theoretical limitations of Vanilla GANs which is caused by the Jensen-Shannon distance. Recall notation (3).